OpenStudy (anonymous):

Consider the line that passes through the point (-1, 6) and has a slope of -4. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

Part one: (y-(-1))= -4(x-6) y+1= -4x+24 part two: y+4x=23 -4x -4x y= -4x+23 Part three: ????? There is no definiton for standard form, in my past and current classes we have only ever had two forms of equations in this type of problem.

OpenStudy (anonymous):

thanks : ) . Wouldnt standard form be 4x+y=23??

OpenStudy (anonymous):

The Standard Form equation of a line has the following formula: Ax + By = C where A ≠0 and B ≠0

OpenStudy (anonymous):

idk i never really was asked for anything but point slope form and slope intercept so im assuming standard for is slope intercept form

OpenStudy (anonymous):

and no problem anytime

hartnn (hartnn):

4x+y=23 is correct standard form